In a study of 369,732 cell phone​ users, it was found that 96 developed cancer of the brain or nervous system. Assuming that cell phones have no​ effect, there is a 0.000323 probability of a person developing cancer of the brain or nervous system. We therefore expect about 120 cases of such cancer in a group of 369,732 people. Estimate the probability of 96 or fewer cases of such cancer in a group of 369,732 people. What do these results suggest about media reports that cell phones cause cancer of the brain or nervous ​system?

a) P(x ≤ 96​) =
Answer #1

mean = 120
sd = sqrt(p * (1-p) * n)
sd = sqrt(0.000323 * (1 - 0.000323) * 369732)
sd = 10.9263

a)
P(X <= 96)
= P(X < 96.5)
= P(z < (96.5 - 120)/10.9263)
= P(z < -2.1508)
= 0.0157